OpenStudy (zned6559):

The ability to roll your tongue is a dominant trait that Alfred Sturtevant observed in 72% of people in a population study. What is the allele frequency of the tongue rolling allele?

OpenStudy (aaronq):

Disclaimer: i haven't done this in a while so i could be wrong, but i think this is how it's done. Since it's dominant, both, homozygous dominant and heterozygous individuals will carry the allele. (\(q\) here is the dominant allele and \(p\) is the recessive) \(q^2+2pq=72\%\) we also know that \(p^2=100\%-72\%\) so \(p=\sqrt{.28}=0.5292\) back to the first part: \(q^2+2(0.5292)q=0.72\) \(q=0.4708\) the frequency of the dominant allele is 0.471